News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Saturday 23 September 2000 12.00pm EDT

During the last week, Chandra completed the observing schedule as planned.

The observatory passed through the seventh and eighth eclipses of the Fall 2000 season with nominal power and thermal performance. A short eclipse software patch was uplinked on 16 Sept in support of the final (8th) eclipse. A software patch to revert to the default medium eclipse state was uplinked on 18 Sept.

A real-time ACIS squeegee mode test was conducted at 2:40pm EDT on 15 Sept to test the squeegee approach with a full raw mode on the S0 chip. Initial assessment of the real-time data shows performance as expected.

The LETG will be inserted during a real-time pass at 1am EDT on 25 Sept for the upcoming week of LETG observations.

The planned observations for the coming week are as follows.

------------------------------------------
Abell 4059              ACIS-S       Sep 24 
NGC 1360                HRC-S/LETG   Sep 25 
Radiation Belts                      Sep 26 
Sirius B                HRC-S/LETG
X Persei (HD 24534)     HRC-S/LETG   Sep 27 
H1504+65                HRC-S/LETG
RX J0019.8+2156         HRC-S/LETG   Sep 28 
Radiation Belts
YY Gem                  HRC-S/LETG   Sep 29 
AM Her                  HRC-S/LETG   Sep 30 
NGC 3516                ACIS-S/LETG
V1494 Aql               HRC-S/LETG   Oct 1  
Radiation Belt
Z eta Pup               HRC-S/LETG   Oct 2

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports